2 Price Discrimination

A monopolist produces a good with constant marginal cost equal to c with c < 1. Assume for now that all consumers have the demand Q(p) = 1 − p. The population is of size 1.

(a) Suppose that the monopolist cannot discriminate in any way among the consumers and has to charge a uniform price, pU . Calculate the price that maximizes profits and the profits that correspond to this price. (10 marks)

(b) Suppose now that the monopolist can charge a two-part tariff (m, p) where m is the fixed fee and p is the price per unit. Expenditure for each con- sumers then is m + p · q. Calculate the two-part tariff that maximizes profits and the profits that correspond to this tariff. (10 marks)

(c) Compare pU and p and comment briefly. Calculate and then compare the situation with a uniform price and a two-part tariff in terms of social welfare. (10 marks)

(d) Assume now instead that there are two types of consumers. The consumers of type 1 have the demand Q1(p) = 1 − p, and the consumers of type 2 have the demand Q2(p) = 1 − p/2. The population is of size 1 and there are equally many consumers of the two types. Calculate the two-part tariff that maximizes the profits of the monopolist. (10 marks)

(e) Compare the two-part tariffs found in questions (b) and (d) and comment briefly. Calculate and compare the two situations in term of social welfare. (10 marks) 



1
Expert's answer
2021-02-19T11:26:34-0500

"(a)q=1-p"

"P=1-Q"

"Mr=1-2q"

"Mc=1"

"Mr=mc"

"1-2Q=1"

"Q=0"

"Pu=1"

"Profit=0"

(B) "1-q=1"

"Q=0"

"P=1"

"M=2"

(C) price increases when a fixed fee is charged.

"(D)p=1-q ,p=1-q\/2 ,mc=1"

"1-q=1 ,q=0"

"1-q\/2=1 ,q=0"

Low income will pay "2"

High income will pay "3"

(E) All consumers will pay "2" when there is no price discrimination .high income will pay "3" when there is price discrimination.